# Hierarchy of timescales in a disordered spin-$1/2$ XX ladder

### v2

@@ IV.2 Thouless times from diffusion constants
Figure 6: Time-dependent disorder-averaged diffusion constants $\overline{D^{(\mu)}}{(t)}$ of a spin-$1/2$ XX ladder as a function of time $t$ for disorder strengths $W/J_{\parallel}=0.5$ in (a) and $W/J_{\parallel}=1.5$ in (b) and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. The color of each line corresponds to a specific number of rungs, $L$. Solid lines represent the spin-diffusion constant $\overline{D^{(\mathrm{S})}}{(t)}$, while dashed lines denote the energy-diffusion constant $\overline{D^{(\mathrm{E})}}{(t)}$. Each diffusion constant is averaged over $M=1000$ disorder realizations. The time evolution is performed using the Lanczos algorithm with a discrete time step of $\Delta t=0.01/J_{\parallel}$ and a maximum of 30 Lanczos iterations.
We plot the time-dependent disorder-averaged diffusion constant of spin transport (solid lines) and energy transport (dashed lines), indicated by $\overline{D^{(\mathrm{S})}}{(t)}$ and $\overline{D^{(\mathrm{E})}}{(t)}$, as a function of time at the coupling ratio $J_{\perp}/J_{\parallel}=1.0$, and the disorder strengths $W/J_{\parallel}=0.5$ in Fig. 6(a) and $W/J_{\parallel}=1.5$ in Fig. 6(b). The line colors represent four different numbers of rungs, with $L\in\{7,8,9,10\}$.
At weak disorder strengths, exemplified by $W/J_{\parallel}=0.5$ in Fig. 6(a), the disorder-averaged diffusion constant $\overline{D^{(\mu)}}{(t)}$ exhibits an initial increase, driven by the positive current autocorrelation function $\overline{C^{(\mu)}}{(t)}>0$ at short times. However, in contrast to the clean limit [80], this growth is transient, and $\overline{C^{(\mu)}}{(t)}$ eventually undergoes a first sign change at a certain time (followed by oscillations and more sign changes). For stronger disorder strengths, such as $W/J_{\parallel}=1.5$ in Fig. 6(b), the sign change in $\overline{C^{(\mu)}}{(t)}$ occurs at an earlier time, marking a more rapid departure from the initial growth regime.
The slow approach towards a constant long-time value in Fig. 6 is caused by a transient, presumably subdiffusive behavior. For sufficiently large system sizes, we expect this transient behavior to give way to standard diffusion as seen in other disordered models [134, 135, 136, 133, 67, 137], leading to the formation of a plateau in the time-dependent diffusion constant at late times and thus providing an approximate value for the diffusion constant for our case. We also note that increasing the disorder strength accelerates the formation of the late-time plateau for small system sizes, as shown in Fig. 6.
@@ IV.3.1 Disordered spin-$1/2$ XX ladder
Figure 7: Relaxation timescales of a spin-$1/2$ XX ladder as a function of the number of the rungs, $L$, at disorder strengths $W/J_{\parallel}=1.25$ in (a) and $W/J_{\parallel}=1.5$ in (b) and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. The blue circles with error bars denote the RMT times $t_{\mathrm{RMT}}$ derived from the SFF. The blue dashed lines correspond to the fits of the function $y(L)=a_{0}L^{2}$ to the RMT time $t_{\mathrm{RMT}}$. The spin-Thouless time $t^{(\mathrm{S})}_{\mathrm{Th}}$ (purple lines with shaded bands) and the energy-Thouless time $t^{(\mathrm{E})}_{\mathrm{Th}}$ (green lines with shaded bands) are obtained via the combined mean and standard deviation of the disorder-averaged diffusion constant $\overline{D^{(\mu)}}{(t)}$ computed over the interval $200\leq tJ_{\parallel}\leq 250$ for $L=9,10$. The uncertainty of the diffusion constants is within the linewidth. The number of disorder realizations is $M=5000$ for $t_{\mathrm{RMT}}$ and $M=1000$ for $t_{\mathrm{Th}}$. The other parameters used to obtain this figure can be found in Figs. 5 and 6.
Building on our previous results from Figs. 5 and 6, we finally show the relaxation timescales as a function of the number of rungs, $L$, for a coupling ratio of $J_{\perp}/J_{\parallel}=1.0$ and two different disorder strengths: $W/J_{\parallel}=1.25$ in Fig. 7(a) and $W/J_{\parallel}=1.5$ in Fig. 7(b). The blue circles with error bars represent the RMT times $t_{\mathrm{RMT}}$ from Eq. (8), which are obtained using the Heisenberg time $t_{\mathrm{H}}$. To estimate $t_{\mathrm{RMT}}$ at larger values of $L$, we also fit the function $y(L)=a_{0}L^{2}$ to the data for $t_{\mathrm{RMT}}$ (dashed lines) for the set $L\in\{5,6,7,8\}$. Likewise, the solid lines with shaded bands indicate the Thouless times $t^{(\mu)}_{\mathrm{Th}}$ defined in Eq. (9). The energy-Thouless time $t^{(\mathrm{E})}_{\mathrm{Th}}$ is represented by the green color, while the spin-Thouless time $t^{(\mathrm{S})}_{\mathrm{Th}}$ is denoted by the purple color. As mentioned in Subsec. III.2, the computation of $t^{(\mu)}_{\mathrm{Th}}$ involves estimating the time-independent disorder-averaged diffusion constants $\overline{D^{(\mu)}}$, by considering a portion of the late-time behavior at sufficiently large values of $L$, which yields their combined mean value (solid lines) and standard deviation (shaded bands). The width of those bands is smaller than the linewidth.
@@
As was previously noted in Fig. 6, and also evident from Fig. 7, spin diffusion faster processes than energy diffusion in the spin-$1/2$ XX ladder. This is reflected in
@@
We also note that $t^{(\mathrm{S})}_{\mathrm{Th}}<t^{(\mathrm{E})}_{\mathrm{Th}}$ is seen at all disorder strength including the clean model. This hierarchy of Thouless times shown in Fig. 7 for a spin‑$1/2$ XX ladder is different from other examples of non‑integrable models studied previously [138, 139, 137].
IV.3.2 Disordered spin-$1/2$ XXZ chain with next-nearest neighbor couplings
Figure 8: Relaxation timescales of a spin-$1/2$ XXZ chain with next-nearest neighbor couplings as a function of system size, $L$, at disorder strengths $W/J=0.75$ in (a) and $W/J=1.0$ in (b), nearest-neighboring anisotropy $\Delta/J=1.0$, and next-nearest-neighboring anisotropy $\Delta_{2}/J=0.3$. The blue circles with error bars denote the RMT times $t_{\mathrm{RMT}}$ derived from the SFF. The blue dashed lines correspond to the fits of the function $y(L)=a_{0}L^{2}$ to the RMT time $t_{\mathrm{RMT}}$. The spin-Thouless time $t^{(\mathrm{S})}_{\mathrm{Th}}$ (purple lines with shaded bands) and the energy-Thouless time $t^{(\mathrm{E})}_{\mathrm{Th}}$ (green lines with shaded bands) are obtained via the combined mean and standard deviation of the disorder-averaged diffusion constant $\overline{D^{(\mu)}}{(t)}$ computed over the interval $200\leq tJ\leq 250$ for $L=18,20$. The uncertainty of the diffusion constants is within the linewidth. The number of disorder realizations is $M=5000$ for $t_{\mathrm{RMT}}$ and $M=1000$ for $t_{\mathrm{Th}}$. The other parameters used to obtain this figure can be found in Figs. 5 and 6.
In fact, slower energy diffusion relative to spin diffusion is indeed an exception because there are other non‑integrable models that do not exhibit this unusual hierarchy. To check that, we consider a disordered spin-$1/2$ XXZ chain with next-nearest neighbor couplings. In the literature, it has been established that this model obeys ETH in its clean limit ($W=0$) [108, 107], while its transport properties were studied in [79, 109, 69, 106].
@@ IV.3.2 Disordered spin-$1/2$ XXZ chain with next-nearest neighbor couplings
The first component, $H_{0}$, describes the spin couplings along the chain and is given by
$H_{0}=\frac{J}{2}\sum_{\ell=1}^{L}\quantity(S^{+}_{\ell}S^{-}_{\ell+1}+S^{+}_{\ell}S^{-}_{\ell+1})\,$, (33)
where $J$ represents the coupling strength between neighboring spins, and $L$ denotes the length of the chain. The second component, $H_{\Delta}$, accounts for the longitudinal coupling between nearest-neighbor spins and is expressed as
@@
with $\Delta$ representing the nearest-neighbor anisotropy. The third component, $H_{\Delta_{2}}$, is similar to $H_{\Delta}$ but describes the coupling between next-nearest-neighbor spins, given by
@@ IV.4 Discussion
Let us discuss the key results: (i) the fact that the RMT time is larger than all Thouless times and (ii) that in the spin-1/2 XX ladder, energy transport is slower than spin transport. Our results are consistent with the usual picture that the $L^{2}$ dependence of the RMT time is inherited from diffusive transport. The fact that the RMT is larger than the largest Thouless time may have several reasons. First of all, the SFF measures the entire spectrum and therefore, it sees contributions from the temperature dependence of the diffusion constant, while we computed the diffusion constants at infinite temperature from the Kubo formalism. At sufficiently large systems, the contributions to $t_{\mathrm{RMT}}$ should be dominated by the bulk of the spectrum, though. Technically, there is no unique way of extracting the RMT time from the SFF as we detailed before, and hence $t_{\mathrm{RMT}}$ is only determined up to a prefactor. The difference $t_{\mathrm{RMT}}-t_{\mathrm{Th}}^{(\mathrm{E})}$ is, however, larger than this uncertainty according to our analysis. Conceptually, the prefactor between $t_{\mathrm{RMT}}$ and the largest Thouless time may simply not be one, as one could check by a calculation of the diffusion constant from the spectrum with twisted boundary conditions [67], which is left for future work. Finally, it cannot be completely ruled out that another unidentified relaxation timescale is larger than the Thouless times and closer $t_{\mathrm{RMT}}$, even though this appears unlikely in this model.
@@
for the spin ladder. In order to understand this behavior, it is instructive to consider the clean limit. Special to XX models, there are no diagonal terms at all in the computational basis (i.e., the joint eigenbasis of all $S^{z}_{\ell}$ operators). Since we compute the diffusion constants at infinite temperature, we can use any basis. In the computational basis, it is obvious that a spin excitation can move between different rungs (as the center of the local energy density) without changing the local energy, since the energy expectation value vanishes in all basis states. Consequently, we expect spin to be able to move faster than energy at infinite temperature in the absence of diagonal terms.
To substantiate this picture, we also carried out simulations for a spin-$1/2$ XXZ ladder with $\Delta>0$ (results not shown here). At sufficiently large values of $\Delta$, we find $D^{(\mathrm{E})}>D^{(\mathrm{S})}$, consistent with the results of [139] obtained for Heisenberg ladders. Whether a finite nonzero value of the exchange anisotropy $\Delta$ is required to reverse the order of the diffusion constants or not is left for future work.
In principle, disorder introduces diagonal terms, coupling the modification of the local spin moment to the local energy density again. Our results suggest that, at a sufficiently large disorder strength $W$, the diffusion constants converge because the local energy density picks up a term proportional to the local spin density, hence tying energy transport to spin transport. An interesting open question is the connection of our results to the existence of ballistic and localized subspaces in certain spin ladder models [79, 105], which we leave for future work.
@@ V Conclusions
We investigated the non-equilibrium dynamics of a disordered spin-$1/2$ XX ladder by utilizing a pair of diagnostic measures from quantum chaos and transport. Specifically, we employed the spectral form factor to extract the timescale associated with the onset of random matrix theory dynamics, i.e., the RMT time. Additionally, we used the infinite-temperature diffusion constant to characterize the system’s different transport properties and to determine the Thouless times associated with spin and energy diffusion.
We found that the RMT time scales quadratically with linear-system size suggesting that it is controlled by diffusive transport. Moreover, we found that the RMT time is larger than the Thouless times. We pointed out several possible reasons for this quantitative difference.
Among the two transport channels in the spin-$1/2$ XX ladder, the spin-Thouless time is smaller than the energy-Thouless time. We argue that this is due to the absence of any diagonal terms in the computational basis, implying that spin can move without affecting the local energy density. We expect that most spin models should exhibit the more generic behavior, i.e., $D^{(\mathrm{E})}>D^{(\mathrm{S})}$. We verified this for a non-integrable spin-$1/2$ XXZ chain and for spin-1/2 XXZ ladders, where the RMT time again scales with $L^{2}$ yet $t_{\mathrm{Th}}^{(\mathrm{E})}<t_{\mathrm{Th}}^{(\mathrm{S})}$.

### v3

@@ IV.2 Thouless times from diffusion constants
Figure 6: Time-dependent disorder-averaged diffusion constants $\overline{D^{(\mu)}}{(t)}$ of a spin-$1/2$ XX ladder as a function of time $t$ for disorder strengths (a) $W/J_{\parallel}=0.5$ and (b) $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. The color of each line corresponds to a specific number of rungs $L$. Solid lines represent the spin-diffusion constant $\overline{D^{(\mathrm{S})}}{(t)}$, while dashed lines denote the energy-diffusion constant $\overline{D^{(\mathrm{E})}}{(t)}$. Each diffusion constant is averaged over $M=1000$ disorder realizations. The time evolution is performed using the Lanczos algorithm with a discrete time step of $\Delta t=0.01/J_{\parallel}$ and a maximum of 30 Lanczos iterations.
We plot the time-dependent disorder-averaged diffusion constant of spin transport (solid lines) and energy transport (dashed lines), indicated by $\overline{D^{(\mathrm{S})}}{(t)}$ and $\overline{D^{(\mathrm{E})}}{(t)}$, as a function of time at the coupling ratio $J_{\perp}/J_{\parallel}=1.0$ and the disorder strengths $W/J_{\parallel}=0.5$ in Fig. 6(a) and $W/J_{\parallel}=1.5$ in Fig. 6(b). The line colors represent four different numbers of rungs, with $L\in\{7,8,9,10\}$.
At weak disorder strengths, exemplified by $W/J_{\parallel}=0.5$ in Fig. 6(a), the disorder-averaged diffusion constant $\overline{D^{(\mu)}}{(t)}$ exhibits an initial increase, driven by the positive current autocorrelation function $\overline{C^{(\mu)}}{(t)}>0$ at short times. However, in contrast with the clean limit [79], this growth is transient, and $\overline{C^{(\mu)}}{(t)}$ eventually undergoes a first sign change at a certain time (followed by oscillations and more sign changes). For stronger disorder strengths, such as $W/J_{\parallel}=1.5$ in Fig. 6(b), the sign change in $\overline{C^{(\mu)}}{(t)}$ occurs at an earlier time, marking a more rapid departure from the initial growth regime.
The slow approach toward a constant long-time value in Fig. 6 is caused by a transient, presumably subdiffusive behavior. For sufficiently large system sizes, we expect this transient behavior to give way to standard diffusion as seen in other disordered models [136, 137, 138, 135, 66, 107], leading to the formation of a plateau in the time-dependent diffusion constant at late times and thus providing an approximate value for the diffusion constant for our case. We also note that increasing the disorder strength accelerates the formation of the late-time plateau for small system sizes, as shown in Fig. 6.
@@ IV.3.1 Disordered spin-$1/2$ XX ladder
Figure 7: Relaxation timescales of a spin-$1/2$ XX ladder as a function of the number of the rungs, $L$, at disorder strengths (a) $W/J_{\parallel}=1.25$ and (b) $W/J_{\parallel}=1.5$ and coupling ratio $J_{\perp}/J_{\parallel}=1.0$. The blue circles with error bars denote the RMT times $t_{\mathrm{RMT}}$ derived from the SFF, and these error bars represent the systematic uncertainties in extracting $t_{\mathrm{RMT}}$. The blue dashed lines correspond to the fits of the function $y(L)=a_{0}L^{2}$ to the RMT time $t_{\mathrm{RMT}}$. The spin-Thouless time $t^{(\mathrm{S})}_{\mathrm{Th}}$ (purple lines with shaded bands) and the energy-Thouless time $t^{(\mathrm{E})}_{\mathrm{Th}}$ (green lines with shaded bands) are obtained via the combined mean and standard deviation of the disorder-averaged diffusion constant $\overline{D^{(\mu)}}{(t)}$ computed over the interval $200\leq tJ_{\parallel}\leq 250$ for $L=9,10$. The uncertainty of the diffusion constants is within the linewidth. The number of disorder realizations is $M=5000$ for $t_{\mathrm{RMT}}$ and $M=1000$ for $t_{\mathrm{Th}}$. The other parameters used to obtain this figure can be found in Figs. 5 and 6.
Building on our previous results from Figs. 5 and 6, we finally show the relaxation timescales as a function of the number of rungs $L$ for a coupling ratio of $J_{\perp}/J_{\parallel}=1.0$ and two different disorder strengths: $W/J_{\parallel}=1.25$ in Fig. 7(a) and $W/J_{\parallel}=1.5$ in Fig. 7(b). The blue circles with error bars represent the RMT times $t_{\mathrm{RMT}}$ from Eq. (8), which are obtained using the Heisenberg time $t_{\mathrm{H}}$. To estimate $t_{\mathrm{RMT}}$ at larger values of $L$, we also fit the function $y(L)=a_{0}L^{2}$ to the data for $t_{\mathrm{RMT}}$ (dashed lines) for the set $L\in\{5,6,7,8\}$. We do not extract the exponent independently; therefore, our procedure only demonstrates a consistent $L^{2}$ scaling. Likewise, the solid lines with shaded bands indicate the Thouless times $t^{(\mu)}_{\mathrm{Th}}$ defined in Eq. (9). The energy-Thouless time $t^{(\mathrm{E})}_{\mathrm{Th}}$ is represented by the green color, while the spin-Thouless time $t^{(\mathrm{S})}_{\mathrm{Th}}$ is denoted by the purple color. As mentioned in Sec. III.2, the computation of $t^{(\mu)}_{\mathrm{Th}}$ involves estimating the time-independent disorder-averaged diffusion constants $\overline{D^{(\mu)}}$, by considering a portion of the late-time behavior at sufficiently large values of $L$, which yields their combined mean value (solid lines) and standard deviation (shaded bands). The width of those bands is smaller than the linewidth.
@@
As was previously noted in Fig. 6 and evident from Fig. 7, spin diffusion processes faster than energy diffusion in the spin-$1/2$ XX ladder. This is reflected in
@@
We also note that $t^{(\mathrm{S})}_{\mathrm{Th}}<t^{(\mathrm{E})}_{\mathrm{Th}}$ is seen at all disorder strengths investigated, i.e., $W/J_{\parallel}\lesssim 2$ including the clean model. This hierarchy of Thouless times shown in Fig. 7 for a spin‑$1/2$ XX ladder is different from other examples of nonintegrable models studied previously [105, 106, 107].
IV.3.2 Disordered spin-$1/2$ XXZ chain with next-nearest-neighbor couplings
Figure 8: Relaxation timescales of a spin-$1/2$ XXZ chain with next-nearest-neighbor couplings as a function of system size $L$ at disorder strengths (a) $W/J=0.75$ and (b) $W/J=1.0$, nearest-neighboring anisotropy $\Delta/J=1.0$, and next-nearest-neighboring anisotropy $\Delta_{2}/J=0.3$. The blue circles with error bars denote the RMT times $t_{\mathrm{RMT}}$ derived from the SFF, and these error bars represent the systematic uncertainties in extracting $t_{\mathrm{RMT}}$. The blue dashed lines correspond to the fits of the function $y(L)=a_{0}L^{2}$ to the RMT time $t_{\mathrm{RMT}}$. The spin-Thouless time $t^{(\mathrm{S})}_{\mathrm{Th}}$ (purple lines with shaded bands) and the energy-Thouless time $t^{(\mathrm{E})}_{\mathrm{Th}}$ (green lines with shaded bands) are obtained via the combined mean and standard deviation of the disorder-averaged diffusion constant $\overline{D^{(\mu)}}{(t)}$ computed over the interval $200\leq tJ\leq 250$ for $L=18,20$. The uncertainty of the diffusion constants is within the linewidth. The number of disorder realizations is $M=5000$ for $t_{\mathrm{RMT}}$ and $M=1000$ for $t_{\mathrm{Th}}$. The other parameters used to obtain this figure can be found in Figs. 5 and 6.
In fact, slower energy diffusion relative to spin diffusion is indeed an exception because there are other nonintegrable models that do not exhibit this unusual hierarchy. To check that, we consider a disordered spin-$1/2$ XXZ chain with next-nearest-neighbor couplings. In the literature, it has been established that this model obeys ETH in its clean limit ($W=0$) [110, 109], while its transport properties were studied in Refs. [78, 111, 68, 108].
@@ IV.3.2 Disordered spin-$1/2$ XXZ chain with next-nearest-neighbor couplings
The first component $H_{0}$ describes the spin couplings along the chain and is given by
[…]
where $J$ represents the coupling strength between neighboring spins, and $L$ denotes the length of the chain. The second component $H_{\Delta}$ accounts for the longitudinal coupling between nearest-neighbor spins and is expressed as
@@
with $\Delta$ representing the nearest-neighbor anisotropy. The third component $H_{\Delta_{2}}$ is similar to $H_{\Delta}$ but describes the coupling between next-nearest-neighbor spins, given by
@@ IV.4 Discussion
Let us discuss the key results: (i) the fact that the RMT time is larger than all Thouless times and (ii) that in the spin-1/2 XX ladder, energy transport is slower than spin transport. Our results are consistent with the usual picture that the $L^{2}$ dependence of the RMT time is inherited from diffusive transport. The fact that the RMT is larger than the largest Thouless time may have several reasons. First of all, the SFF measures the entire spectrum and therefore, it sees contributions from the temperature dependence of the diffusion constant, while we computed the diffusion constants at infinite temperature from the Kubo formalism. At sufficiently large systems, the contributions to $t_{\mathrm{RMT}}$ should be dominated by the bulk of the spectrum, though. Technically, there is no unique way of extracting the RMT time from the SFF as we detailed before, and hence, $t_{\mathrm{RMT}}$ is only determined up to a prefactor. The difference $t_{\mathrm{RMT}}-t_{\mathrm{Th}}^{(\mathrm{E})}$ is, however, larger than this uncertainty according to our analysis. Conceptually, the prefactor between $t_{\mathrm{RMT}}$ and the largest Thouless time may simply not be one, as one could check by a calculation of the diffusion constant from the spectrum with twisted boundary conditions [66], which is left for future work. Finally, it cannot be completely ruled out that another unidentified relaxation timescale is larger than the Thouless times and closer $t_{\mathrm{RMT}}$, even though this appears unlikely in this model.
@@
for the spin ladder. To understand this behavior, it is instructive to consider the clean limit. Special to XX models, there are no diagonal terms at all in the computational basis (i.e., the joint eigenbasis of all $S^{z}_{\ell}$ operators). Since we compute the diffusion constants at infinite temperature, we can use any basis. In the computational basis, it is obvious that a spin excitation can move between different rungs (as the center of the local energy density) without changing the local energy since the energy expectation value vanishes in all basis states. Consequently, we expect spin to be able to move faster than energy at infinite temperature in the absence of diagonal terms.
To substantiate this picture, we also carried out simulations for a spin-$1/2$ XXZ ladder with $\Delta>0$ (results not shown here). At sufficiently large values of $\Delta$, we find $D^{(\mathrm{E})}>D^{(\mathrm{S})}$, consistent with the results of Ref. [106] obtained for Heisenberg ladders. Whether a finite nonzero value of the exchange anisotropy $\Delta$ is required to reverse the order of the diffusion constants is left for future work.
In principle, disorder introduces diagonal terms, coupling the modification of the local spin moment to the local energy density again. Our results suggest that, at a sufficiently large disorder strength $W$, the diffusion constants converge to the same value because the local energy density picks up a term proportional to the local spin density, hence tying energy transport to spin transport. An interesting open question is the connection of our results to the existence of ballistic and localized subspaces in certain spin-ladder models [78, 104], which we leave for future work.
@@ V Conclusions
We investigated the nonequilibrium dynamics of a disordered spin-$1/2$ XX ladder by utilizing a pair of diagnostic measures from quantum chaos and transport. Specifically, we employed the SFF to extract the timescale associated with the onset of RMT, i.e., the RMT time. Additionally, we used the infinite-temperature diffusion constant to characterize the different transport properties of the system and to determine the Thouless times associated with spin and energy diffusion.
We found that the RMT time scales quadratically with linear system size, suggesting that it is controlled by diffusive transport. Moreover, we found that the RMT time is larger than the Thouless times. We pointed out several possible reasons for this quantitative difference.
Among the two transport channels in the spin-$1/2$ XX ladder, the spin-Thouless time is smaller than the energy-Thouless time. We argue that this is due to the absence of any diagonal terms in the computational basis, implying that spin can move without affecting the local energy density. We expect that most spin models should exhibit the more generic behavior, i.e., $D^{(\mathrm{E})}>D^{(\mathrm{S})}$. We verified this for a nonintegrable spin-$1/2$ XXZ chain and for spin-1/2 XXZ ladders, where the RMT time again scales with $L^{2}$ yet $t_{\mathrm{Th}}^{(\mathrm{E})}<t_{\mathrm{Th}}^{(\mathrm{S})}$.